OpenStudy (anonymous):

In a right triangle, 1/5 of the length of the longer leg is equal to 7/12 of the length of the shorter leg. What is the ratio of the length of the hypotenuse to the length of the longer leg?

OpenStudy (anonymous):

37/35

OpenStudy (anonymous):

and can u explain how u got that?

OpenStudy (anonymous):

step by step

OpenStudy (anonymous):

(1/5)L = (7/12)S Therefore: L = (5)(7/12)S = (35/12)S S = S L = (35/12)S Pythagorean Theorem: a^2 + b^2 = c^2 \[c=\sqrt(a ^{2}+b ^{2})\]

OpenStudy (anonymous):

\[Hypotenuse = \sqrt{S^{2}+L ^{2}}\]

OpenStudy (anonymous):

\[H = \sqrt{(S)^{2}+((35/12)S)^{2}} = {37 \over 12}S\]

OpenStudy (anonymous):

We want the ratio of Hypotenuse:LongerSide

OpenStudy (anonymous):

H:L = \[H:L = {37 \over 12}S:{35 \over 12}S\]

OpenStudy (anonymous):

The S's and 12's cancel each other out.

OpenStudy (anonymous):

Leaving us with: 37:35 or 37/35

OpenStudy (anonymous):

Make sense?

OpenStudy (anonymous):

Don't forget to mark a Best Response. Thanks and take care.